# Benchmark for Antibody Binding Affinity Maturation and Design

### v1

@@ 3.3 Evaluation Tasks
3.3.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log-Likelihood
@@
To measure how well a model’s zero-shot predicted log-likelihood aligns with wet-lab verified affinity, we calculated the Spearman correlation between the model likelihood and experimentally measured binding affinity. A high correlation indicates that the model assigns a higher likelihood to strong binders, suggesting that it can identify affinity-enhancing mutations in a zero-shot setting. To evaluate how effectively a model prioritizes the most promising antibodies, we also reported 5-fold precision@10 – the proportion of top 10 ranked variants that achieve at least 5-fold improvement in binding affinity compared to the wild type. The calculation of affinity fold change is detailed in Supplement LABEL:sup:sec:computationalMetrics. We harmonized the log-likelihood computation for all models under a unified setting: the input unit is the mutant-antigen complex, and the output is the likelihood of that complex. The details of likelihood computation across different types of models are provided in Supplement LABEL:calculation_details. In addition to the model’s log likelihood, we report two biophysics-based affinity score as a baseline: binding free energy ($\Delta G$) and the relative solvent-accessible surface areas ($SASA$) of epitope residues (Sec. 3.3.2, Supplement LABEL:sup:sec:computationalMetrics). Lower values of both metrics imply stronger binding. To ensure consistent directionality with model log-likelihoods, we report $-\Delta G$ and $-SASA$.
3.3.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
@@
To assess whether protein models can generate antibody variants with improved binding to a specific antigen, we conducted a case study using F045-092 [56], a naturally occurring antibody that targets the hemagglutinin (HA) protein of influenza H3N2. Notably, F045-092 fails to bind the H1N1 strain California2009 due to a steric barrier at the HA receptor-binding site [16, 64]. No experimentally determined structure of the F045-H1N1 complex exists in public databases, and this particular pair has not been studied in prior antibody design literature. As such, our study represent a completely novel setting, free from data leakage. The structure used in our study was computationally predicted using AlphaFold3. We used four representative models—ESM-IF, SaProt, DiffAb, and MEAN—to redesign the CDR-H3 loop of F045-092 while allowing up to five substitution to maintain H3N2 cross-reactivity. Each model generated 1,500 CDR-H3 variants from an input consisting of the masked F045-092 sequence and a predicted complex structure with the H1N1 HA1 protein (Supplement LABEL:sup:sec:case_study). Sampling strategies varied by model, including autoregressive prediction (SaProt), greedy heuristics (ESM-IF), and diffusion-based or graph-based generation (DiffAb, MEAN; see Supplement LABEL:sup:sec:sec:design_cdrh3 for details).
Figure 3: Evaluation of new antibody design. Left: assess sequence plausibility and binding energy change. Right: high-confidence ranking based on AF3-predicted complex structures, evaluating complex structure confidence, epitope SASA, and inverse folding likelihood.
@@
We evaluated the generated antibody variants from two perspectives: sequence plausibility and binding potential. Sequence plausibility was assessed using the log-likelihood from AntiBERTy [61], reflecting how closely mutations align with natural antibody evolution. We also computed inverse folding likelihoods using ProteinMPNN to determine whether each sequence is compatible with its backbone structure—higher scores indicate greater foldability. Binding potential was evaluated using three structure-based metrics: (1) binding free energy ($\Delta G$) to estimate Ab–Ag interaction strength, (2) epitope SASA to measure differences in the solvent-accessible antigen surface area upon binding, and (3) complex structure confidence (AlphaFold pLDDT [2]) as a proxy for interface stability (Supplement LABEL:sup:sec:computationalMetrics).
@@
Some metrics require the structure of the antibody variant in complex with antigen. Predicting the complex structure of all sampled variants using a folding model is computationally expensive. To address this, we used a two-phased screening approach (Fig. 3). In Phase 1 we evaluated all 1,500 variants per model using sequence plausibility (AntiBERTy likelihood) and $\Delta G$ – metrics that do not require structure prediction – and selected top 20% antibody variants. In Phase 2, we generated full Ab-Ag complex structures with AlphaFold 3 for this subset and computed pLDDT, epitope SASA, and inverse folding likelihood. We identified the final candidates as those in the Pareto-optimal set across all five metrics. We also examined the diversity of the selected antibody designs. We compared antibody PLM embeddings (AntiBERTy), sequence similarity (cdrDist), and structural deviation of the CDR-H3 loop (cdrRMSD) relative to the wild type (Supplement LABEL:sup:sec:computationalMetrics).
@@ 4 Results
4.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log-Likelihood
@@
Figure 5: Proportion of top-10 ranked antibody designs achieving $\geq$5-fold affinity improvement across models and datasets. Only datasets reporting affinity as $-\!\log K_{d}$ were used. Datasets based on enrichment scores were excluded, as enrichment reflects relative sequence abundance and cannot determine fold change. Models marked with * are structure-informed.
As a result, inverse folding models achieved the highest accuracy in predicting experimental binding affinity, attaining the highest average Spearman correlation (Fig. 4) and highest 5-fold precision@10 (Fig. 5) for all inverse folding models tested. This high accuracy may stem from the model’s broader structural scope. PLMs with local structure token have consistently performed best in general protein binding tasks [41, 67], but the same did not hold for antibodies. Inverse-folding models, which encode the entire Ab–Ag complex as a single global representation, consistently outperform models that rely on local structural tokens—such as SaProt, ProSST, and ESM-3. Besides, purely autoregressive sequence models without structure like ProGen2 and ProtGPT2 likewise fail to achieve competitive accuracy on the affinity prediction task. We shuffled the chain order of autoregressive models, but it did not increase the accuracy either (Table LABEL:tab:autoregressive, Supplement LABEL:sup:sec:sec:chain_order). From a structural biology perspective, the inverse folding strategy is effective because protein function is ultimately determined by its three-dimensional structure, which is encoded by the underlying sequence. Mutations that maintain or improve structural integrity have a higher potential to enhance functional properties [63]. On the other hand, by conditioning on the full antibody-antigen structure, inverse folding models can capture long-range residue interactions and contextual features at the binding interface—both of which are critical for affinity but often missed by models relying only on local sequence information [59, 73]. Moreover, among non–inverse folding models, SaProt which leverages local structural representations achieves the second-best performance, particularly in the 5-fold precision@10 metric.
Interestingly, we also found that antibody-finetuned PLMs have varying impacts on zero-shot correlation to binding affinity. When we compare general PLM ESM-2 to its Ab-finetuned model CurrAb [5], CurrAb improved Spearman correlation by +0.074 across all datasets and precision@10 by +0.067. However, AntiFold [26], which is an ESM-IF finetuned on structure of antibodies (OAS [57]) and Ab-Ag complex (SAbDab [15]) data, deteriorated affinity correlation by -0.097 and precision@10 by -0.078 across all datasets, suggesting limited gains on Ab-specific finetuning and potential catastrophic forgetting. [^3] Note that AntiFold reports higher Spearman correlation with 1mlc dataset (0.427) – see Fig S9 [26]). This discrepancy occurs because AntiFold is evaluated using light and heavy chain variants measuring log-likelihood of the CDR regions, whereas AbBiBench excludes light chain mutants and report log-likelihood values of the entire Ab-Ag complex.
Figure 6: A. Binding potential and sequence plausibility in Phase 1. Binding potential was presented as negative binding energy changes $-\Delta\Delta G$, and $\Delta\Delta G$ is defined as $\Delta G_{variant}-\Delta G_{wild}$, where $\Delta G_{wild}=66.34$. The top 20% of variants, defined by plausibility and binding affinity, are shown in the upper right corner. B. Top 20 % variants’ CDR-H3 sequence diversity by a UMAP plot of sequence embeddings. 18 non-dominated Pareto-optimal variants, which are marked with a $\bm{\times}$ symbol, were identified by considering all five metrics on binding potential and sequence plausibility.
Figure 7: Binding potential and sequence plausibility in Phase 2. A. Sequence plausibility measured by ProteinMPNN likelihood. B. Binding potential measured by epitope SASA (presented as negative change, $-\Delta SASA$) . C. Binding potential measured by pLDDT of Ab-Ag complex structure.
@@ 4.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
We evaluated whether selected models could generate antibody variants with stronger binding affinity to H1N1 than the wild-type F045-092. Each model produced 1,500 CDR-H3 variants with up to five mutations. As DiffAb often generated wild-type sequences with altered structures, we excluded duplicates, resulting in 467 valid DiffAb variants and a total of 4,967 unique variants across all models. In Phase 1 screening, ESM-IF and SaProt produced variants with both strong binding potential and high sequence plausibility. Their average $\Delta\Delta G$ values were -29.27 and -20.79, respectively, indicating substantial improvement in binding energy (Fig. 6, LABEL:fig:phase1A). These variants also retained AntiBERTy plausibility scores (-0.663 for ESM-IF and -0.666 for SaProt) close to that of the wild type (-0.655; Fig. LABEL:fig:phase1B). On the other hands, DiffAb generated plausible sequences (-0.657) but failed to improve binding energy ($\Delta\Delta G=2.03$), suggesting mutations without affinity gain. In contrast, MEAN produced variants with improved binding ($\Delta\Delta G=-13.95$), but at the cost of lower sequence plausibility (-0.680), indicating a trade-off between biophysical fitness and evolutionary realism (Fig. LABEL:fig:phase1).
Figure 8: Relationship between structure confidence (pLDDT) and structural deviation from wild type (cdrRMSD) of CDR-H3 generated by models.
@@
Among the top 20% of Phase 1 variants, 158, 91, and 26 candidates were selected from ESM-IF, SaProt, and MEAN, respectively. In Phase 2, SaProt achieved the highest sequence plausibility (ProteinMPNN log-likelihood = -1.406) and complex structure integrity (pLDDT = 83.22), followed by ESM-IF (-1.415, pLDDT = 82.82; Fig.  7LABEL:fig:phase2-2).
@@
To assess the sequence diversity, we visualized CDR-H3 embeddings using AntiBERTy (Fig. 6B). MEAN variants formed a tight, isolated cluster, while ESM-IF and SaProt occupied broader, partially overlapping regions—indicating greater diversity. Structural comparison revealed that SaProt generated CDR-H3 variants closest to wild type (mean cdrRMSD = 2.28), followed by ESM-IF (2.70) and MEAN (3.25). SaProt also produced variants with higher CDR-H3 structural confidence (pLDDT), suggesting better foldability (Fig. 8LABEL:fig:diversity-cdrh3). Sequence divergence analysis using cdrDist showed SaProt variants were most similar to wild type (0.222), compared to ESM-IF (0.279) and MEAN (0.277) (Fig. 8, LABEL:fig:diversity-sw). This aligns with the average number of substitutions per CDR-H3: 2.98 for SaProt, 3.03 for MEAN, and 3.75 for ESM-IF.
@@ 5 Conclusion and Limitations
This study introduces AbBiBench, a biologically relevant and structurally informed benchmarking framework for antibody binding affinity maturation and optimization. Recognizing the limitations of traditional computational evaluation metrics, our approach explicitly incorporates antibody-antigen complex information, thereby aligning more closely with the biological realities of antibody interactions. Our results demonstrate that global structure-informed protein language model used for inverse folding methods, such as ESM-IF and ProteinMPNN, outperform other evaluated computational models, primarily due to their effective integration of structural context. In a case study focused on redesigning the F045-092 antibody for binding to the H1N1 influenza subtype, we identified 18 Pareto-optimal antibody variants with improved predicted affinity and structural integrity. These 18 variants have been successfully expressed in vitro, and wet lab experiments to quantify their binding affinity to H1N1 hemagglutinin are currently underway. The results of these assays will directly validate the computational predictions and help assess the true binding potential of the designed antibodies. Limitations of this study include the lack of currently available experimental neutralization data and the relatively small size of some benchmark datasets (Supplement LABEL:sup:sec:limitations). In future work, we plan to expand the benchmark to include additional therapeutic properties such as stability, immunogenicity, and developability, and to incorporate functional assay data (e.g., IC50) to further align model evaluations with biological outcomes.
@@
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the contributions, inclusion of 11 antibody-antigen datasets, and the evaluation of protein models for affinity prediction and antibody design. These contributions are supported by empirical results in Sections 3 and 4.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: A dedicated discussion of limitations is provided in Section 5 and Supplement Section LABEL:sup:sec:limitations, including the lack of experimental neutralization assays and limited dataset size for some antigens.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper does not present any formal theoretical results, proofs, or theorems.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper?
  15. Answer: [Yes]
  16. Justification: The paper includes detailed descriptions of datasets (Section 3.1), model settings, evaluation tasks (Section 3.3), and provides full dataset/code access via Hugging Face and GitHub (Supplement Section LABEL:sup:resources).
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results?
  19. Answer: [Yes]
  20. Justification: Dataset is publicly available at Hugging Face (https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset) and code is available via GitHub (https://github.com/MSBMI-SAFE/AbBiBench).
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details?
  23. Answer: [Yes]
  24. Justification: Sections 3.23.3 and Supplement Materials Section LABEL:calculation_detailsLABEL:sup:sec:case_study include all necessary details on model selection, sampling strategies, and evaluation metrics.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. 8.
    Experiments compute resources
  29. Question: Does the paper provide sufficient information on the computer resources?
  30. Answer: [Yes]
  31. Justification: As illustrated in Supplement Materials LABEL:sup:resources, model training is not required in this study. All inference tasks are conducted on a single NVIDIA H100 80GB GPU per model.
  32. 9.
    Code of ethics
  33. Question: Does the research conducted in the paper conform with the NeurIPS Code of Ethics?
  34. Answer: [Yes]
  35. Justification: The work follows ethical standards in open-source data sharing, does not involve human subjects or dual-use risks, and conforms with NeurIPS ethical principles.
  36. 10.
    Broader impacts
  37. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  38. Answer: [Yes]
  39. Justification: The manuscript highlights implications for therapeutic antibody development and model misuse risks are low given the biological specificity of antibody design (Section 5 and case study).
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards for responsible release of high-risk data/models?
  42. Answer: [N/A]
  43. Justification: The dataset and models do not pose high risk for misuse, as they are focused on antibody sequence design for therapeutic research.
  44. 12.
    Licenses for existing assets
  45. Question: Are existing assets (code, data, models) properly credited and licensed?
  46. Answer: [Yes]
  47. Justification: All reused datasets and models are properly cited (e.g., FLAb, ProteinMPNN, SaProt, etc.) with sources listed in sections 2, 3.2 and Supplement Materials LABEL:sup:data_details.
  48. 13.
    New assets
  49. Question: Are new assets well documented and is documentation provided?
  50. Answer: [Yes]
  51. Justification: The AbBiBench dataset is introduced with detailed documentation on Hugging Face and Croissant format metadata (see Supplement Materials LABEL:sup:resources).
  52. 14.
    Crowdsourcing and research with human subjects
  53. Question: For crowdsourcing experiments and research with human subjects, does the paper include full instructions and compensation?
  54. Answer: [N/A]
  55. Justification: The work does not involve any human participants or crowdsourcing.
  56. 15.
    Institutional review board (IRB) approvals
  57. Question: Does the paper describe IRB approval for human subject research?
  58. Answer: [N/A]
  59. Justification: No research involving human subjects was conducted.
  60. 16.
    Declaration of LLM usage
  61. Question: Does the paper describe LLM usage if it is an original or non-standard component?
  62. Answer: [N/A]
  63. Justification: This study evaluates pre-trained protein language models (PLMs), which are specialized models trained on protein sequences, not general-purpose large language models (LLMs). No LLMs were used in developing or implementing the core methods of this research.

### v2

@@ 3.3 Evaluation Tasks
3.3.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log‑Likelihood
@@
To measure how well a model’s zero-shot predicted log-likelihood aligns with wet-lab verified affinity, we calculated the Spearman correlation between the model likelihood and experimentally measured binding affinity. A high correlation indicates that the model assigns a higher likelihood to strong binders, suggesting that it can identify affinity-enhancing mutations in a zero-shot setting. To evaluate how effectively a model prioritizes the most promising antibodies, we also reported 5-fold precision@10 – the proportion of top 10 ranked variants that achieve at least 5-fold improvement in binding affinity compared to the wild type. The calculation of affinity fold change is detailed in Supplement LABEL:sup:sec:computationalMetrics. We harmonized the log-likelihood computation for all models under a unified setting: the input unit is the mutant-antigen complex, and the output is the likelihood of that complex. The details of likelihood computation across different types of models are provided in Supplement LABEL:calculation_details. In addition to the model’s log likelihood, we report two biophysics-based affinity score as a baseline: binding free energy ($\Delta G$) and the relative solvent-accessible surface areas ($SASA$) of epitope residues (Sec. 3.3.2, Supplement LABEL:sup:sec:computationalMetrics). Lower values of both metrics imply stronger binding. To ensure consistent directionality with model log-likelihoods, we report $\Delta G$ and SASA-SASA.
3.3.2 Generate Antibody Variants with Strong H1N1 Influenza Affinity
@@
To assess whether protein models can generate antibody variants with improved binding to a specific antigen, we conducted a case study using F045-092 Ohshima et al. (2011), a naturally occurring antibody that targets the hemagglutinin (HA) protein of influenza H3N2. Notably, F045-092 fails to bind the H1N1 strain California2009 due to a steric barrier at the HA receptor-binding site Ekiert et al. (2012); Simmons et al. (2023). No experimentally determined structure of the F045-H1N1 complex exists in public databases, and this particular pair has not been studied in prior antibody design literature. As such, our study represents a completely novel setting, free from data leakage. The structure used in our study was computationally predicted using AlphaFold3. We used four representative models—ESM-IF, SaProt, DiffAb, and MEAN—to redesign the CDR-H3 loop of F045-092 while allowing up to five substitutions to maintain H3N2 cross-reactivity. We focus on generating mutations in the variable heavy chain (vH) due to its high diversity from V(D)J recombination and the central role of CDR-H3 in antigen binding Chen et al. (2024). The vH often acts as a unique antigen-specific signature Davies & Riechmann (1995), while the light chain often remains relatively conserved across functional antibodies Jaffe et al. (2022), making vH the most relevant region for affinity optimization. Each model generated 1,500 CDR-H3 variants from an input consisting of the masked F045-092 sequence and a predicted complex structure with the H1N1 HA1 protein (Supplement LABEL:sup:sec:case_study). Sampling strategies varied by model, including autoregressive prediction (SaProt), greedy heuristics (ESM-IF), and diffusion-based or graph-based generation (DiffAb, MEAN; see Supplement LABEL:sup:sec:sec:design_cdrh3 for details).
@@
We evaluated the generated antibody variants from two perspectives: sequence plausibility and binding potential. Sequence plausibility was assessed using the log-likelihood from AntiBERTy Ruffolo et al. (2021), reflecting how closely mutations align with natural antibody evolution. We also computed inverse folding likelihoods using ProteinMPNN to determine whether each sequence is compatible with its backbone structure—higher scores indicate greater foldability. Binding potential was evaluated using three structure-based metrics: (1) binding free energy ($\Delta G$) to estimate Ab–Ag interaction strength, (2) epitope SASA to measure differences in the solvent-accessible antigen surface area upon binding, and (3) complex structure confidence (AlphaFold pLDDT Abramson et al. (2024)) as a proxy for interface stability (Supplement LABEL:sup:sec:computationalMetrics).
@@
To identify the final candidates, we used a two-phased screening approach (Fig. LABEL:fig:evaluating). In Phase 1 we evaluated all 1,500 variants per model using sequence plausibility (AntiBERTy likelihood) and $\Delta G$ – metrics that do not require structure prediction – and selected top 20% antibody variants. In Phase 2, we generated full Ab-Ag complex structures with AlphaFold 3 for this subset and computed pLDDT, epitope SASA, and inverse folding likelihood. We identified the final candidates as those in the Pareto-optimal set across all five metrics. We also examined the diversity of the selected antibody designs. We compared antibody PLM embeddings (AntiBERTy), sequence similarity (cdrDist), and structural deviation of the CDR-H3 loop (cdrRMSD) relative to the wild type (Supplement LABEL:sup:sec:computationalMetrics).
@@ 4 Results
4.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log‑Likelihood
As a result, inverse folding models achieved the highest accuracy in predicting experimental binding affinity, attaining the highest average Spearman correlation (Fig. 3) and highest 5-fold precision@10 (Fig. 4) for all inverse folding models tested. This high accuracy may stem from the model’s broader structural scope. PLMs with local structure token have consistently performed best in general protein binding tasks Li et al. (2024); Su et al. (2023), but the same did not hold for antibodies. Inverse-folding models, which encode the entire Ab–Ag complex as a single global representation, consistently outperform models that rely on local structural tokens—such as SaProt, ProSST, and ESM-3. Besides, purely autoregressive sequence models without structure like ProGen2 and ProtGPT2 likewise fail to achieve competitive accuracy on the affinity prediction task. We shuffled the chain order of autoregressive models, but it did not increase the accuracy either (Table LABEL:tab:autoregressive, Supplement LABEL:sup:sec:sec:chain_order). From a structural biology perspective, the inverse folding strategy is effective because protein function is ultimately determined by its three-dimensional structure, which is encoded by the underlying sequence. Mutations that maintain or improve structural integrity have a higher potential to enhance functional properties Shanker et al. (2024). On the other hand, by conditioning on the full antibody-antigen structure, inverse folding models can capture long-range residue interactions and contextual features at the binding interface—both of which are critical for affinity but often missed by models relying only on local sequence information Orlandi et al. (2020); Wang et al. (2018).
@@ 4.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log‑Likelihood
Figure 4: Proportion of top-10 ranked antibody designs achieving $\geq$5-fold affinity improvement across models and datasets. Only datasets reporting affinity as $\log K_{d}$ were used. Datasets based on enrichment scores were excluded, as enrichment reflects relative sequence abundance and cannot determine fold change. Models marked with * are structure-informed.
Moreover, among non–inverse folding models, SaProt which leverages local structural representations achieves the second-best performance, particularly in the 5-fold precision@10 metric. Interestingly, we also found that antibody-finetuned PLMs have varying impacts on zero-shot correlation to binding affinity. When we compare general PLM ESM-2 to its Ab-finetuned model CurrAb Burbach & Briney (2025), CurrAb improved Spearman correlation by +0.074 across all datasets and precision@10 by +0.067. However, AntiFold Høie et al. (2025), which is an ESM-IF finetuned on structure of antibodies (OAS Olsen et al. (2022a)) and Ab-Ag complex (SAbDab Dunbar et al. (2014)) data, deteriorated affinity correlation by -0.097 and precision@10 by -0.078 across all datasets, suggesting limited gains on Ab-specific finetuning and potential catastrophic forgetting. [^3] Note that AntiFold reports higher Spearman correlation with 1mlc dataset (0.427) – see Fig S9 Høie et al. (2025)). This discrepancy occurs because AntiFold is evaluated using light and heavy chain variants measuring log-likelihood of the CDR regions, whereas AbBiBench excludes light chain mutants and report log-likelihood values of the entire Ab-Ag complex.
Figure 5: Binding potential and sequence plausibility in Phase 2. A. Sequence plausibility measured by ProteinMPNN likelihood. B. Binding potential measured by epitope SASA (presented as negative change, $\Delta SASA$) . C. Binding potential measured by pLDDT of Ab-Ag complex structure.
@@ 4.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
We evaluated whether selected models could generate antibody variants with stronger binding affinity to H1N1 than the wild-type F045-092. Each model produced 1,500 CDR-H3 variants with up to five mutations. As DiffAb often generated wild-type sequences with altered structures, we excluded duplicates, resulting in 467 valid DiffAb variants and a total of 4,967 unique variants across all models. In Phase 1 screening, ESM-IF and SaProt produced variants with both strong binding potential and high sequence plausibility. Their average $\Delta\Delta G$ values were -29.27 and -20.79, respectively, indicating substantial improvement in binding energy (Fig. LABEL:fig:phase1-screening, LABEL:fig:phase1A). These variants also retained AntiBERTy plausibility scores (-0.663 for ESM-IF and -0.666 for SaProt) close to that of the wild type (-0.655; Fig. LABEL:fig:phase1B). On the other hands, DiffAb generated plausible sequences (-0.657) but failed to improve binding energy ($\Delta\Delta G=2.03$), suggesting mutations without affinity gain. In contrast, MEAN produced variants with improved binding ($\Delta\Delta G=-13.95$), but at the cost of lower sequence plausibility (-0.680), indicating a trade-off between biophysical fitness and evolutionary realism (Fig. LABEL:fig:phase1).
Figure 6: ELISA $\text{OD}_{450}$ signals for 21 model-designed F045-092 mutants and controls against H1 hemagglutinin. Three antibodies were used as controls, including the wild-type F045-092 (H3-specific binder), 5J8 (H1-specific binder), and FI6v3 (a broadly neutralizing antibody for both subtypes). Diamonds represent mutants with higher $\text{OD}_{450}$ values than F045-092 against H3.
@@
Among the top 20% of Phase 1 variants, 158, 91, and 26 candidates were selected from ESM-IF, SaProt, and MEAN, respectively. In Phase 2, SaProt achieved the highest sequence plausibility (ProteinMPNN log-likelihood = -1.406) and complex structure integrity (pLDDT = 83.22), followed by ESM-IF (-1.415, pLDDT = 82.82; Fig.  5LABEL:fig:phase2-2).
To assess the sequence diversity, we visualized CDR-H3 embeddings using AntiBERTy (Fig. LABEL:fig:phase1-screeningB). MEAN variants formed a tight, isolated cluster, while ESM-IF and SaProt occupied broader, partially overlapping regions—indicating greater diversity.
@@
Structural comparison revealed that SaProt generated CDR-H3 variants closest to wild type (mean cdrRMSD = 2.28), followed by ESM-IF (2.70) and MEAN (3.25). SaProt also produced variants with higher CDR-H3 structural confidence (pLDDT), suggesting better foldability (Fig. LABEL:fig:diversity-rmsdLABEL:fig:diversity-cdrh3). Sequence divergence analysis using cdrDist showed SaProt variants were most similar to wild type (0.222), compared to ESM-IF (0.279) and MEAN (0.277) (Fig. LABEL:fig:diversity-rmsd, LABEL:fig:diversity-sw). This aligns with the average number of substitutions per CDR-H3: 2.98 for SaProt, 3.03 for MEAN, and 3.75 for ESM-IF.
In vitro experiments.
We performed ELISA binding assays on the 21 designed mutants against hemagglutinin (HA) from A/California/2009 (H1). The Figure 6 shows that all variants produced signals above the detection threshold for H1, indicating a gain of H1 binding activity. This suggests that our in-silico evaluation pipeline enables efficient screening of affinity-enhancing mutations from multiple models, thereby reducing the time and cost associated with experimental validation. Notably, the greatest gain in affinity against the H1 subtype was achieved by the ESM-IF derived designs supporting the conclusion of our in silico evaluation.
@@ 5 Conclusion and Limitations
This study introduces AbBiBench, a biologically relevant and structurally informed benchmarking framework for antibody binding affinity maturation and optimization. Recognizing the limitations of traditional computational evaluation metrics, our approach explicitly incorporates antibody-antigen complex information, thereby aligning more closely with the biological realities of antibody interactions. Our results demonstrate that global structure-informed protein language model used for inverse folding methods, such as ESM-IF and ProteinMPNN, outperform other evaluated computational models, primarily due to their effective integration of structural context. In a case study focused on redesigning the F045-092 antibody for binding to the H1N1 influenza subtype, we identified 21 Pareto-optimal antibody variants with improved predicted affinity and structural integrity. These 21 variants have been successfully expressed in vitro, and conduct ELISA assays to quantify their binding affinity to H1N1 hemagglutinin. The results of these assays directly validate the computational predictions and help assess the true binding potential of the designed antibodies. Limitations of this study include the lack of currently available experimental neutralization data and the relatively small size of some benchmark datasets (Supplement LABEL:sup:sec:limitations). In future work, we plan to expand the benchmark to include additional therapeutic properties such as stability, immunogenicity, and developability, and to incorporate functional assay data (e.g., IC50) to further align model evaluations with biological outcomes.